# LATex: Leveraging Attribute-based Text Knowledge for Aerial-Ground Person Re-Identification

### v2

@@ IV-B Implementation Details
Our proposed framework is implemented with PyTorch on one NVIDIA A100 GPU. We use the pre-trained CLIP-Base-16 as our backbone. All input images are resized to $256\times 128$. To enhance the generalization ability, we utilize multiple augmentation techniques such as random horizontal flipping, padding and random erasing [66] for all inputs. While the model is training, the mini-batch size is 128 consisted of 16 identities and 8 instances of each identity. We fine-tune the model with Adam optimizer [67] with a base learning rate of 3.5e-4. A learning rate scheduling strategy is employed, combining a warm-up phase with cosine decay and a scaling factor of 0.01. The total epoch is 120. For the hyper-parameters, we set $\lambda_{1}$ and $\lambda_{2}$ in Eq. 11 to 0.25 and 1.0, respectively.
LATex†. LATex† is a variant of LATex that adopts the full fine-tuning strategy to ensure a fair comparison with other full fine-tuning AG-ReID methods. More specifically, we unfreeze all trainable parameters of the pre-trained CLIP vision and text encoders and update them with a learning rate of 5e-6. Other experimental settings, including the learning rate decay strategy, optimizer, and the number of training epochs, are kept consistent with LATex.
@@
Metric ViT VDT LATex
Trainable Params(M) 86.24 85.90 35.97
GPU Memory(G) 0.075 0.078 0.079
Inference Speed(s/batch) 0.35 0.41 0.67
Flops(G) 11.37 11.46 15.33
@@ IV-C Performance Comparison
On the AG-ReID.v1, our LATex achieves a Rank-1 accuracy of 88.88% and an mAP of 79.19% under the evaluation protocol G→A. Compared with SeCap, our LATex represents improvements of 1.87% in Rank-1 and and 0.85% in mAP, respectively. As for the evaluation protocol A→G, the performance of our LATex is 84.41% Rank-1 and 75.85% mAP, showing very competitive results. The consistent improvements on two evaluation protocols clearly demonstrate the importance of leveraging attribute information in AG-ReID. On two large-scale datasets, CARGO and AG-ReID.v2, our LATex achieves highly competitive performance. It is worth noting that, since there are no attribute annotations on CARGO, we remove PACG. As a result, the final outputs of AIE, i.e., $\tilde{P}^{L}$, are directly passed to CPT as attribute representations. Our LATex achieves superior performance with fewer trainable parameters, fully validating its effectiveness.
@@ IV-D Training and Inference Cost Comparison
In Tab. IV, we provide a comprehensive comparison of the training and inference costs on AG-ReID.v1. During training, our model significantly reduces the trainable parameters to about 40% compared with VDT. As a result, when all other conditions remain the same, the resource requirements for training the model (e.g., the computational overhead of parameter update) can be reduced. As for the inference cost, LATex obtains notable performance gains in the cost of introducing minimal computational overhead compared with other AG-ReID methods. The inference speed is also very comparable.
@@ IV-E Ablation Studies
To demonstrate the effectiveness of our proposed modules, we evaluate them on the AG-ReID.v1 dataset. The results are shown in Tab. V. Furthermore, we conduct comprehensive and evaluations to explore the details of our model design.
@@
Effect of Trainable Prompts. Fig. 3 shows the effect of the number of learnable prompts $P^{i}_{v}$. With the increase of trainable parameters, the performance of LATex remains stable, demonstrating its robustness to this factor.
@@
Effect of View Token. Tab. IX shows the effectiveness of view token. Even though the variant without view token demonstrates competitive performance, explicitly incorporating view token helps LATex further enhance its cross-view retrieval capabilities. Since view token is camera-specific and can be pre-defined, we integrate this token into the CPT to achieve the optimal performance.

### v3

@@ IV-B Implementation Details
Our proposed framework is implemented with PyTorch on one NVIDIA A100 GPU. We use the pre-trained CLIP-Base-16 as our backbone. All input images are resized to $256\times 128$. To enhance the generalization ability, we utilize multiple augmentation techniques such as random horizontal flipping, padding and random erasing [56] for all inputs. While the model is training, the mini-batch size is 128 consisted of 16 identities and 8 instances of each identity. We fine-tune the model with Adam optimizer [57] with a base learning rate of 3.5e-4. A learning rate scheduling strategy is employed, combining a warm-up phase with the cosine decay and a scaling factor of 0.01. The total epoch is 120. For the hyper-parameters, we set $\lambda_{1}$ and $\lambda_{2}$ in Eq. 11 to 0.25 and 1.0, respectively. The end-to-end training process takes 1.5 hours.
LATex†. LATex† is a variant of LATex that adopts the full fine-tuning strategy to ensure a fair comparison with other full fine-tuning AG-ReID methods. Specifically, we unfreeze all trainable parameters of the pre-trained CLIP’s vision and text encoders and update them with a learning rate of 5e-6. Other settings, including the learning rate decay strategy, optimizer, and the number of training epochs, are kept consistent with LATex. The end-to-end training process takes about 2.5 hours.
@@
Metric ViT VDT LATex LATex†
Trainable Params(M) 86.24 85.90 35.97 122.00
GPU Memory(G) 0.075 0.078 0.079 0.108
Inference Speed(s/batch) 0.35 0.41 0.67 0.67
Flops(G) 11.37 11.46 15.33 15.33
@@ IV-C Performance Comparison
On the AG-ReID.v1, our LATex achieves a Rank-1 accuracy of 88.88% and an mAP of 79.19% under the evaluation protocol G→A. Compared with SeCap, our LATex presents improvements of 1.87% in Rank-1 and and 0.85% in mAP, respectively. As for the evaluation protocol A→G, the performance of our LATex is 84.41% Rank-1 and 75.85% mAP, showing very competitive results. The consistent improvements on two evaluation protocols clearly demonstrate the importance of leveraging attribute information in AG-ReID. On two large-scale datasets, CARGO and AG-ReID.v2, our LATex achieves highly competitive performance. It is worth noting that, there are no attribute annotations on CARGO. To address the lack of attribute annotations, we remove PACG and directly use the output visual prompts from AIE as pseudo-attribute representations for CPT. This adaptation ensures that our method can still leverage structural text prompts even in the absence of explicit attribute labels. As a result, it can be used for attribute-missing benchmarks, such as CARGO. The superior performance on CARGO fully validates its effectiveness in attribute-sparse domains.
@@ IV-D Training and Inference Cost Comparison
In Tab. IV, we provide a comprehensive comparison of the training and inference costs on AG-ReID.v1. Our proposed LATex significantly reduces trainable parameters compared with full fine-tuning methods like VDT and our LATex†. Although LATex† achieves the highest performance, it requires substantially more trainable parameters and higher GPU memory usage. In contrast, LATex offers a more efficient solution with competitive results. Notably, the inference costs are determined by the model architecture. Thus, the inference speed and FLOPs are the same between LATex and LATex†.
@@ IV-E Ablation Studies
To demonstrate the effectiveness of our proposed modules, we evaluate them on the AG-ReID.v1 dataset. The results are shown in Tab. V. Furthermore, we conduct comprehensive and evaluations to investigate the details of our model design.
@@
Effect of Trainable Prompts. Fig. 3 shows the effect of the number of attribute prompts. With the increase of trainable parameters, the performance of LATex remains stable, demonstrating its robustness to this factor.
@@
Effect of View Tokens. Tab. IX shows the effectiveness of view tokens. Even though the variant without view token demonstrates competitive performance, explicitly incorporating view token s helps LATex further enhance its cross-view retrieval capabilities. Since view tokens are camera-specific and can be pre-defined, we integrate these tokens into the CPT to achieve the optimal performance.
Performance Upper Bound Analysis. To explore the performance upper bound, we remove PACG and directly incorporate attribute labels into the [attribute tokens] placeholder of CPT. It simulates a scenario with perfect attribute predictions. The last row of Tab. IX shows that our method achieves notably high performance with ground truth attributes. These results demonstrates a strong theoretical upper bound of our framework and its potential for further optimization.